# SAGe: A Lightweight Algorithm-Architecture Co-Design for Alleviating Data Preparation Overheads in Large-Scale Genome Analysis

### v1

@@ Abstract.
Genome analysis plays an important role in many critical fields. Therefore, there have been extensive efforts to accelerate genome analysis by alleviating the challenges of analyzing exponentially growing volumes of genomic data. Prior works on accelerating genome analysis typically assume that the data is ready to be analyzed in the desired format; in real usage scenarios, however, it is common practice to store genomic data in storage systems in a compressed format. Unfortunately, preparing genomic data (i.e., accessing the compressed data from the storage system, decompressing the read data, and reformatting it) for a genome analysis accelerator leads to large performance and energy overheads and, thus, significantly diminishes the potential intended benefits of the acceleration efforts. To actually harness the full benefits of the extensive efforts on accelerating genome analysis, without resorting to storing massive volumes of genomic data uncompressed, there is a critical need to effectively address data preparation overheads. The solution must meet three criteria: (ihigh performance and energy efficiency for data preparation, (iihigh compression ratios, comparable to state-of-the-art genomic compression algorithms, and (iii) be lightweight to enable seamless integration with a broad range of genome analysis systems, regardless of where they are implemented. Simultaneously optimizing for all of these criteria is challenging, particularly due to the high decompression complexity of state-of-the-art genomics-specific compression algorithms (which, despite their high compression ratios, lead to large performance/energy overheads and require costly hardware resources) and resource constraints of a wide range of genome analysis systems.
We propose SAGe, an algorithm-system co-design that enables highly-compressed storage and high-performance access of large-scale genomic data in the desired format. Driven by our rigorous analysis of genomic datasets’ features, we propose an efficient and synergistic co-design of new (de)compression algorithm, lightweight hardware units, storage data layout, and interface commands. To achieve high performance, energy efficiency, and low hardware cost when preparing data, SAGe encodes data in lightweight structures that can be interpreted and decoded by efficient sequential scans and lightweight hardware. To maintain high compression ratios using only these lightweight structures, SAGe exploits unique features of genomic data to enhance space efficiency. SAGe’s lightweight design can be seamlessly integrated with a broad range of existing acceleration techniques for genome analysis to, for the first time, enable unleashing the performance and energy benefits of genome analysis acceleration, without relying on large-scale genomic data to be stored in uncompressed forms. Our evaluation shows that SAGe improves the average end-to-end performance and energy efficiency of state-of-the-art genome-analysis accelerators by 3.0$\times$–12.3$\times$ and 18.8$\times$–49.6$\times$, respectively, compared to when the accelerators rely on state-of-the-art decompression tools.
@@ 1. Introduction
Genome analysis, which refers to the analysis of the genomic information encoded in an organism’s DNA, plays an important role in many critical fields, such as personalized medicine (clark2019diagnosis, ; farnaes2018rapid, ; sweeney2021rapid, ; alkan2009personalized, ; flores2013p4, ; ginsburg2009genomic, ; chin2011cancer, ; Ashley2016, ), tracing outbreaks of communicable diseases (bloom2021massively, ; yelagandula2021multiplexed, ; le2013selected, ; nikolayevskyy2016whole, ; qiu2015whole, ; gilchrist2015whole, ), and ensuring food safety through pathogen monitoring (e002244, ; TONG2021130, ). To be able to analyze genomic information computationally, a DNA sample from an organism undergoes a process called sequencing, which converts the information from DNA molecules in the sample to digital data. This digital information is encoded in the form of sequences that represent the DNA nucleotides with the characters A, C, G, and T. Current sequencing technologies cannot sequence long DNA molecules end-to-end. Instead, state-of-the-art sequencers generate randomly- and redundantly-sampled smaller and inexact fragments of these molecules, known as reads. Sets of genomic reads (called read sets) are then used in genome analysis. The importance of genome analysis, along with the rapid improvement in sequencing (i.e., reduced costs and increased throughput (berger2023navigating, )), has led to the rapidly increasing adoption of genome analysis in recent years (clark2019diagnosis, ; farnaes2018rapid, ; sweeney2021rapid, ; ginsburg2009genomic, ; chin2011cancer, ; Ashley2016, ; bloom2021massively, ; gilchrist2015whole, ) and continuous exponential growth in genomic data generation (katz2021sra, ; ena2022, ; srastats, ; enastats, ) (far outpacing Moore’s law (stephens2015big, )).
@@
Due to its importance, there have been extensive efforts to accelerate genome analysis by alleviating the challenges of analyzing exponentially growing volumes of genomic data [^1] We focus on genomic read sets, which are the predominant form of genomic data (enastats, ).. These efforts focus on alleviating computational overheads via algorithmic optimizations (e.g., (zhang2000greedy, ; slater2005automated, ; li2018minimap2, ; myers1999fast, ; marco2021fast, ; marcosola2023optimal, ; grootkoerkamp2024apa2, )) or hardware accelerators (e.g., (mutlu2023accelerating, ; alser2022molecules, ; lou2020helix, ; lou2018brawl, ; shahroodi2023swordfish, ; saavedra2020mining, ; markus2020benchmarking, ; subramaniyan2021accelerated, ; huangfu2018radar, ; khatamifard2021genvom, ; cali2020genasm, ; gupta2019rapid, ; li2021pim, ; angizi2019aligns, ; zokaee2018aligner, ; turakhia2018darwin, ; fujiki2018genax, ; madhavan2014race, ; cheng2018bitmapper2, ; houtgast2018hardware, ; houtgast2017efficient, ; zeni2020logan, ; ahmed2019gasal2, ; nishimura2017accelerating, ; de2016cudalign, ; liu2015gswabe, ; liu2013cudasw++, ; liu2009cudasw++, ; liu2010cudasw++, ; wilton2015arioc, ; goyal2017ultra, ; chen2016spark, ; chen2014accelerating, ; chen2021high, ; fujiki2020seedex, ; banerjee2018asap, ; fei2018fpgasw, ; waidyasooriya2015hardware, ; chen2015novel, ; rucci2018swifold, ; haghi2021fpga, ; li2021pipebsw, ; ham2020genesis, ; ham2021accelerating, ; wu2019fpga, )), and/or reducing data movement overheads via near-data processing (e.g., in main memory (wu2021sieve, ; shahroodi2022krakenonmem, ; shahroodi2022demeter, ; dashcam23micro, ; hanhan2022edam, ; zou2022biohd, ; cali2020genasm, ; huangfu2018radar, ; khatamifard2021genvom, ; gupta2019rapid, ; li2021pim, ; angizi2019aligns, ; zokaee2018aligner, ; Zhang_2023_alignerD, ) or storage system (mansouri2022genstore, ; abakus23taco, ; megis, ; jun2016storage, )).
@@
Prior works on accelerating genome analysis typically assume that the data is ready to be analyzed in the desired format; in real usage scenarios, however, it is common practice to store genomic data in storage systems in a compressed format (berger2023navigating, ; zhu2013high, ; Deorowicz2013, ). This is because storing raw (i.e., uncompressed) data in multiple formats for different accelerators is impractical, given the massive volume of genomic data. In fact, due to the importance of efficiently storing large volumes of genomic data in a space-efficient manner, there exist many compression techniques (e.g., (chandak2018spring, ; chandak2017compression, ; kokot2022colord, ; Meng2023, ; roguski2018fastore, ; dufort2021renano, ; kowalski2019pgrc, ; dufort2020enano, )) specialized for genomic data to achieve higher compression ratios compared to state-of-the-art general-purpose compressors.
Unfortunately, preparing genomic data (i.e., accessing the compressed data from the storage system, decompressing the read data, and reformatting it) for a genome analysis accelerator significantly diminishes the potential intended benefits of genome analysis acceleration efforts. Our motivational analysis (§3) shows that this overhead can significantly increase the end-to-end execution time and energy consumption of genome analysis. In particular, we observe that the decompression overhead can be even more significant than the I/O overhead that occurs when reading larger uncompressed files from storage. Our observation highlights a critical trade-off; while it is possible to avoid decompression time by storing data in an uncompressed way, doing so requires allocating significantly larger storage capacity, which is inefficient, costly, and unsustainable for rapidly increasing volumes of genomic data.
@@
To be able to actually harness the full benefits of the extensive efforts on accelerating genome analysis, without resorting to storing massive volumes of genomic data in uncompressed forms, there is a critical need for a design that effectively addresses data preparation overheads while meeting three key requirements. First, it should achieve high performance and energy efficiency when preparing data. Second, it should maintain high compression ratios comparable to state-of-the-art genomics-specific compression techniques. Third, it must be lightweight to enable high versatility for seamless and unobtrusive integration with a broad range of genome analysis systems, including general-purpose systems, specialized accelerators, and near-data processing architectures.
@@
Simultaneously optimizing for all three requirements (i.e., high data preparation performance and energy efficiency, compression ratios comparable to state-of-the-art genomic compression, and lightweightness for seamless integration) is challenging. Some prior works (e.g., (guo2013gpu, ; qiao2019fpga, ; zhao2017streaming, ; jiang2021exma, ; arram2015fpga, ; wang2018accelerating, ; chen2023efficient, ; leavline2013hardware, )) accelerate certain computational kernels (e.g., BWT (zhao2017streaming, ; qiao2019fpga, ), FM-Index search (jiang2021exma, ; arram2015fpga, ; wang2018accelerating, ), and LZMA (chen2023efficient, ; leavline2013hardware, )) that are widely used in many state-of-the-art genomics-specific compression techniques (e.g., (lan2021genozip, ; Meng2023, ; kowalski2019pgrc, ; chandak2017compression, )). Despite their benefits, these approaches require expensive computational units, large buffers, or large DRAM bandwidth or capacity (e.g., due to many random accesses to large data structures for matching patterns). Therefore, none of these approaches is suitable for seamless integration with genome analysis accelerators, particularly when targeting integration with genome analysis systems in resource-constrained environments. For example, while near data processing (NDP) leads to significant benefits in genome analysis (wu2021sieve, ; shahroodi2022krakenonmem, ; shahroodi2022demeter, ; dashcam23micro, ; hanhan2022edam, ; zou2022biohd, ; cali2020genasm, ; huangfu2018radar, ; khatamifard2021genvom, ; gupta2019rapid, ; li2021pim, ; angizi2019aligns, ; zokaee2018aligner, ; Zhang_2023_alignerD, ; mansouri2022genstore, ; abakus23taco, ; megis, ; jun2016storage, ) by fundamentally alleviating data movement overheads when analyzing large-scale genomic data, such designs are often implemented in resource-constrained environments (e.g., within the memory system or storage device), which makes expensive decompression techniques prohibitive to deploy. Consider NDP accelerators implemented inside the storage system (e.g., (mansouri2022genstore, ; abakus23taco, ; megis, ; jun2016storage, )), which provide large performance and energy benefits for genome analysis by (i) fundamentally alleviating data-movement overhead from the storage system and (ii) reducing the application’s burden from the rest of the system (i.e., computational units and main memory). Without the capability to prepare data for analysis inside the SSD’s resource-constrained environment, all such works would either necessitate data to be stored in an uncompressed way, which is inefficient, or require external decompression, which completely undermines the fundamental benefits of NDP.
@@
Our goal in this work is to improve the performance and energy efficiency of genome analysis by mitigating data preparation overheads, while maintaining compression ratios comparable to state-of-the-art genomics-specific compressors, and ensuring a lightweight design that can be seamlessly integrated with a broad range of genome analysis systems. To this end, we propose SAGe, an algorithm-system co-design that enables highly-compressed storage and high-performance access of large-scale genomic data in the desired format. The key idea of SAGe is to leverage the unique features of genomic data in both the compression mechanism and architecture design. SAGe’s lightweight co-design, for the first time, enables unleashing the performance and energy benefits of genome analysis acceleration, without relying on large-scale genomic data to be stored in uncompressed forms.
@@
Key Mechanism. Driven by our rigorous analysis of genomic datasets’ features, we propose an efficient and synergistic co-design of new (de)compression algorithm, lightweight hardware units, storage data layout, and interface commands. SAGe’s co-design comprises four aspects. First, to achieve high performance, energy efficiency, and low hardware cost when preparing data, SAGe encodes data in lightweight structures that can be decoded by efficient sequential scans and lightweight hardware. To this end, we sequentially encode information regarding reads in a genomic read set in lightweight array structures. We efficiently leverage the long-range similarity common in DNA data to compactly, and yet losslessly, encode genomic reads in these structures. We reduce the size of the array structures by adapting the bit width of each entry. To enable correct decoding of data, we use a guide array that records the bit width used for each corresponding array entry. Second, to maintain high compression ratios using only these lightweight structures, SAGe exploits unique features of genomic data by tuning the encoding of array and guide array structures (e.g., bit count granularities) based on the characteristics of each genomic dataset, which are shaped by the sequencing technology and the dataset’s genetic composition. Third, we design lightweight hardware units, which can be seamlessly integrated with any genome analysis system to efficiently interpret and decode SAGe’s data structures with only simple operations and efficient streaming accesses. Fourth, we design an efficient data layout in the storage system that enables leveraging the SSD’s full bandwidth when accessing genomic data. Fifth, we propose specialized interface commands, which are exposed to genome analysis applications to communicate with SAGe’s hardware units to decompress and convert data to the accelerator’s desired format (e.g., two-bit (cheever1989using, ) or one-hot (Hirst1992, ) encoding).
@@
We evaluate SAGe when preparing data for different genome analysis systems. We compare SAGe against configurations where data decompression is performed via (i) pigz (adler2015pigz, ), a widely-used general-purpose compression tool; (ii) Spring (chandak2018spring, ) and NanoSpring (Meng2023, ), state-of-the-art genomics-specific compression tools for short and long reads, respectively; and (iii) Spring (chandak2018spring, ) and NanoSpring (Meng2023, ) with hardware acceleration. We evaluate SAGe’s impact on the end-to-end performance and energy efficiency of genome analysis when integrated with a state-of-the-art accelerator for read mapping (chen2023gem, ) (a fundamental genome analysis task). We show that SAGe improves performance by an average of 12.3$\times$, 3.9$\times$, and 3.0$\times$, and energy efficiency by 49.6$\times$, 24.6$\times$, and 18.8$\times$ compared to configurations where pigz, (Nano)Spring, and hardware-accelerated (Nano)Spring prepare the data for the read mapping accelerator. We also demonstrate SAGe’s benefits in preparing data in resource-constrained environments and its capability for integration with genome analysis systems, regardless of where they are implemented. To this end, we evaluate SAGe’s integration with GenStore (mansouri2022genstore, ), a state-of-the-art NDP genomic accelerator inside the SSD. SAGe provides 30.7$\times$, 9.9$\times$, and 7.5$\times$ average speedup compared to the configurations with pigz, (Nano)Spring, and hardware-accelerated (Nano)Spring, respectively, with only the very low area cost of 0.7% of the three cores (cortexr4, ) in an SSD controller (samsung860pro, ). SAGe provides all of these benefits while achieving significantly larger compression ratios compared to pigz (2.9$\times$ larger, on average) and comparable compression ratios to the state-of-the-art genomics-specific compressors (average reduction of only 4.6%).
@@
  • We demonstrate that preparing data (i.e., accessing data from the storage system and decompressing and reformatting it) for genome analysis accelerators can significantly diminish their end-to-end potential benefits.
  • We propose SAGe, the first algorithm-system co-design for high-compressed storage and high-performance access of large-scale genomic data, able to seamlessly integrate with a wide range of genome analysis systems. SAGe’s lightweight co-design, for the first time, enables unleashing the performance and energy benefits of genome analysis acceleration, without relying on large-scale genomic data to be stored in uncompressed forms.
  • We enable SAGe’s lightweight design by leveraging the features of genomic data in both compression mechanism and architecture design.
  • We rigorously evaluate SAGe and show that, when integrated with state-of-the-art genome sequence analysis systems, SAGe significantly improves their end-to-end performance and energy efficiency, while incurring only a low hardware cost.

### v2

# SAGe: A Lightweight Algorithm-Architecture Co-Design for Mitigating the Data Preparation Bottleneck in Large-Scale Genome Analysis

@@ Abstract.
Given the exponentially growing volumes of genomic data, there are extensive efforts to accelerate genome analysis. We demonstrate a major bottleneck that greatly limits and diminishes the benefits of state-of-the-art genome analysis accelerators: the data preparation bottleneck, where genomic data is stored in compressed form and needs to be decompressed and formatted first before an accelerator can operate on it. To mitigate this bottleneck, we propose SAGe, an algorithm-architecture co-design for highly-compressed storage and high-performance access of large-scale genomic data. SAGe overcomes the challenges of mitigating the data preparation bottleneck while maintaining high compression ratios (comparable to genomic-specific compression algorithms) at low hardware cost. This is enabled by leveraging key features of genomic datasets to co-design (i) a new (de)compression algorithm, (ii) hardware, (iii) storage data layout, and (iv) interface commands to access storage. SAGe stores data in structures that can be rapidly interpreted and decompressed by efficient streaming accesses and lightweight hardware. To achieve high compression ratios using only these lightweight structures, SAGe exploits unique features of genomic data. We show that SAGe can be seamlessly integrated with a broad range of genome analysis hardware accelerators to mitigate their data preparation bottlenecks. Our results demonstrate that SAGe improves the average end-to-end performance and energy efficiency of two state-of-the-art genome analysis accelerators by 3.0$\times$–32.1$\times$ and 18.8$\times$–49.6$\times$, respectively, compared to when the accelerators rely on state-of-the-art decompression tools.
@@ 1. Introduction
Genome analysis plays an important role in many critical fields, such as personalized medicine (clark2019diagnosis, ; farnaes2018rapid, ; sweeney2021rapid, ; alkan2009personalized, ; flores2013p4, ; ginsburg2009genomic, ; chin2011cancer, ; Ashley2016, ), tracing outbreaks of communicable diseases (bloom2021massively, ; yelagandula2021multiplexed, ; le2013selected, ; nikolayevskyy2016whole, ; qiu2015whole, ; gilchrist2015whole, ), and ensuring food safety through pathogen monitoring (e002244, ; TONG2021130, ). To analyze genomic information computationally, a DNA sample from an organism undergoes a process called sequencing, which converts the information from DNA molecules in the sample to digital data. Current sequencing technologies cannot sequence long DNA molecules end-to-end. Instead, state-of-the-art sequencers generate randomly- and redundantly-sampled smaller and inexact DNA fragments, known as reads. Sets of genomic reads (called read sets) are then used in genome analysis. The importance of genome analysis, along with the rapid improvement in sequencing (i.e., reduced costs and increased throughput (berger2023navigating, )), has led to the rapidly increasing adoption of genome analysis in recent years (clark2019diagnosis, ; farnaes2018rapid, ; sweeney2021rapid, ; ginsburg2009genomic, ; chin2011cancer, ; Ashley2016, ; bloom2021massively, ; gilchrist2015whole, ) and continuous exponential growth in genomic data generation (katz2021sra, ; ena2022, ; srastats, ; enastats, ) (far outpacing Moore’s law (stephens2015big, )).
@@
Due to its importance, there are extensive efforts to accelerate genome analysis by alleviating the challenges of analyzing exponentially growing volumes of genomic data. [^1] In this work, we focus on genomic read sets, which are the predominant form of genomic data (enastats, ). These efforts focus on alleviating computational overheads via algorithmic optimizations (e.g., (zhang2000greedy, ; slater2005automated, ; li2018minimap2, ; myers1999fast, ; marco2021fast, ; marcosola2023optimal, ; grootkoerkamp2024apa2, )) or hardware accelerators (e.g., (mutlu2023accelerating, ; alser2022molecules, ; lou2020helix, ; lou2018brawl, ; shahroodi2023swordfish, ; saavedra2020mining, ; markus2020benchmarking, ; subramaniyan2021accelerated, ; huangfu2018radar, ; khatamifard2021genvom, ; cali2020genasm, ; gupta2019rapid, ; li2021pim, ; angizi2019aligns, ; zokaee2018aligner, ; turakhia2018darwin, ; fujiki2018genax, ; madhavan2014race, ; cheng2018bitmapper2, ; houtgast2018hardware, ; houtgast2017efficient, ; zeni2020logan, ; ahmed2019gasal2, ; nishimura2017accelerating, ; de2016cudalign, ; liu2015gswabe, ; liu2013cudasw++, ; liu2009cudasw++, ; liu2010cudasw++, ; wilton2015arioc, ; goyal2017ultra, ; chen2016spark, ; chen2014accelerating, ; chen2021high, ; fujiki2020seedex, ; banerjee2018asap, ; fei2018fpgasw, ; waidyasooriya2015hardware, ; chen2015novel, ; rucci2018swifold, ; haghi2021fpga, ; li2021pipebsw, ; ham2020genesis, ; ham2021accelerating, ; wu2019fpga, )), and/or reducing data movement overheads via near-data processing (e.g., in main memory (wu2021sieve, ; shahroodi2022krakenonmem, ; shahroodi2022demeter, ; dashcam23micro, ; hanhan2022edam, ; zou2022biohd, ; cali2020genasm, ; huangfu2018radar, ; khatamifard2021genvom, ; gupta2019rapid, ; li2021pim, ; angizi2019aligns, ; zokaee2018aligner, ; Zhang_2023_alignerD, ) or storage system (mansouri2022genstore, ; abakus23taco, ; megis, ; jun2016storage, )).
@@
Unfortunately, prior work on accelerating genome analysis ignores what is needed to prepare genomic data for analysis, especially that genomic data is typically stored compressed and needs to be decompressed before it can be processed. It is common practice to store genomic data in storage systems compressed (berger2023navigating, ; zhu2013high, ; Deorowicz2013, ) because storing raw (i.e., uncompressed) data in is impractical, given the massive volumes of genomic data. In fact, due to the importance of storing genomic data in a space-efficient manner, there exist many compression techniques (e.g., (chandak2018spring, ; chandak2017compression, ; kokot2022colord, ; Meng2023, ; roguski2018fastore, ; dufort2021renano, ; kowalski2019pgrc, ; dufort2020enano, )) specialized for genomic data to achieve higher compression ratios than state-of-the-art general-purpose compressors.
Figure 1. Throughput of different genome analysis systems.
Data Preparation Bottleneck in Genome Analysis. We demonstrate that the benefits of prior works on accelerating genome analysis are greatly diminished due to what we call the data preparation bottleneck, where genomic data is stored in compressed form and needs to be decompressed and formatted first before an accelerator can operate on it. Our motivational analysis (§3) shows that this bottleneck can greatly decrease the end-to-end throughput of modern genome analysis accelerators. For example, Fig. 1 shows that the hardware-accelerated system (assuming data is already in the desired uncompressed format) can potentially lead to a significant throughput improvement over the software baseline (by alleviating its computation and memory bottlenecks). However, in practice, the data preparation bottleneck greatly reduces this benefit. In particular, our motivational analysis shows that decompression overhead can be even more significant than the storage I/O overhead of reading larger uncompressed data. This highlights a trade-off; while it is possible to avoid decompression overhead by storing data uncompressed, doing so requires allocating significantly larger storage capacity, which is inefficient, costly, and unsustainable for rapidly growing genomic data volumes.
@@
Requirements for Mitigating the Data Preparation Bottleneck. To fully benefit from the extensive efforts on accelerating genome analysis, without resorting to storing massive volumes of genomic data uncompressed, there is a critical need for a design that effectively mitigates the data preparation bottleneck while meeting three key requirements. First, it should achieve high performance and energy efficiency when preparing data. Second, it should maintain high compression ratios comparable to state-of-the-art genomics-specific compression techniques. Third, it must be lightweight to enable high versatility for seamless and unobtrusive integration with a broad range of genome analysis systems, including general-purpose systems, specialized accelerators, portable genomics systems, and near-data processing architectures.
@@
Challenges. Simultaneously optimizing for all three requirements is challenging. Some prior works (e.g., (guo2013gpu, ; qiao2019fpga, ; zhao2017streaming, ; jiang2021exma, ; arram2015fpga, ; wang2018accelerating, ; chen2023efficient, ; leavline2013hardware, )) accelerate certain computational kernels that are widely used in many state-of-the-art genomics-specific compressors (e.g., (lan2021genozip, ; Meng2023, ; kowalski2019pgrc, ; chandak2017compression, )). Despite their benefits, these approaches require expensive computational units, large buffers, or large DRAM bandwidth or capacity (e.g., due to many random accesses to large data structures for matching patterns). Therefore, none of these approaches is suitable for seamless integration with genome analysis accelerators, particularly when targeting integration with accelerators in resource-constrained environments, such as various near-data processing (NDP) accelerators (e.g., (wu2021sieve, ; shahroodi2022krakenonmem, ; shahroodi2022demeter, ; dashcam23micro, ; hanhan2022edam, ; zou2022biohd, ; cali2020genasm, ; huangfu2018radar, ; khatamifard2021genvom, ; gupta2019rapid, ; li2021pim, ; angizi2019aligns, ; zokaee2018aligner, ; Zhang_2023_alignerD, ; mansouri2022genstore, ; abakus23taco, ; megis, ; jun2016storage, )) or portable genomics devices (e.g., (SmidgION, ; MinIONMk1CO, ; palatnick2020igenomics, ; Ballard2018, ; Oehler2023, )), which play critical roles in genome analysis.
@@
Our goal in this work is to improve the performance and energy efficiency of genome analysis by mitigating the data preparation bottleneck, while maintaining compression ratios comparable to state-of-the-art genomics-specific compressors, and ensuring a lightweight design. To this end, we propose SAGe, an algorithm-architecture co-design for highly-compressed storage and high-performance access of large-scale genomic data, to mitigate the data preparation bottleneck in genome analysis. The key idea of SAGe is to leverage the unique features of genomic datasets (determined by the sequencing technologies and genetic composition) in both the compression algorithm and architecture design, such that the highly-compressed data can be interpreted by lightweight hardware and rapidly prepared for genome analysis.
@@
Key Mechanism. Leveraging our rigorous analysis of genomic datasets’ features, we propose an efficient and synergistic co-design, which comprises five aspects. First, SAGe stores information regarding reads in a genomic read set in array structures that, during decompression, can be interpreted by efficient accesses and lightweight hardware. Second, to maintain high compression ratios using only these lightweight structures, SAGe’s compression algorithm exploits unique features of genomic data by tuning the encoding of the data structures based on the characteristics of each genomic dataset, which are shaped by the sequencing technology and the dataset’s genetic composition. Third, we design lightweight hardware units to efficiently decompress SAGe’s data structures with only simple operations and efficient streaming accesses. SAGe’s low-cost and lightweight hardware enables seamless integration with a wide range of genome analysis accelerators, without competing for resources (e.g., memory bandwidth and capacity) required by genome analysis tasks. Fourth, we design an efficient data layout that enables leveraging the storage system’s full bandwidth when accessing compressed genomic data. Fifth, we propose specialized interface commands, which are exposed to genomics applications to access storage and communicate with SAGe’s hardware to decompress the data to the desired format.
@@
Key Results. We evaluate the impact of SAGe on the end-to-end performance of two state-of-the-art genome analysis accelerators. We compare SAGe against configurations where data decompression is performed via (i) pigz (adler2015pigz, ), a widely-used general-purpose compressor; (ii) Spring (chandak2018spring, ) and NanoSpring (Meng2023, ), state-of-the-art genomics-specific compressors for short and long reads, respectively; and (iii) Spring (chandak2018spring, ) and NanoSpring (Meng2023, ) with hardware acceleration. We evaluate the end-to-end performance of genome analysis when integrated with a state-of-the-art accelerator for read mapping (chen2023gem, ) (a fundamental genome analysis task). We show that SAGe improves performance by an average of 12.3$\times$, 3.9$\times$, and 3.0$\times$, and energy efficiency by 49.6$\times$, 24.6$\times$, and 18.8$\times$ compared to configurations with pigz, (Nano)Spring, and hardware-accelerated (Nano)Spring. We demonstrate SAGe’s benefits in resource-constrained environments by evaluating SAGe’s integration with GenStore (mansouri2022genstore, ), a state-of-the-art NDP genomic accelerator inside the SSD. SAGe provides 32.1$\times$, 10.4$\times$, and 7.8$\times$ average speedup compared to the configurations with pigz, (Nano)Spring, and hardware-accelerated (Nano)Spring, respectively, with only the very low area cost of 0.7% of the three cores (cortexr4, ) in an SSD controller (samsung860pro, ). SAGe provides all of these benefits while achieving significantly larger compression ratios compared to pigz (2.9$\times$ larger, on average) and comparable compression ratios to the state-of-the-art genomics-specific compressors (with the average reduction of only 4.6%).
@@
  • We demonstrate that the data preparation bottleneck can greatly diminish the potential benefits of genome analysis hardware accelerators.
  • We propose SAGe, an algorithm-system co-design for highly-compressed storage and high-performance access of large-scale genomic data, to mitigate the data preparation bottleneck in genome analysis.
  • We enable SAGe’s lightweight design with the key idea to leverage the unique features of genomic data in both the compression algorithm and architecture design, such that the highly-compressed data can be interpreted by lightweight hardware and rapidly prepared for genome analysis.
  • We rigorously evaluate SAGe and show that, when integrated with state-of-the-art hardware genome sequence analysis accelerators, SAGe significantly improves their end-to-end performance and energy efficiency, while incurring only a low hardware cost.